Find the source current is in the following circuit. j1 19 1: 2 10 1. = 42120°A Ideal transformer
Answer #1

Qn) find Is jis Tuz m in 2 Is ☺ žlice Ville V2 BILE VE 30112 Y to Y lo=42120° A f10 = 4120° A Let vi & V2 are the Votage af.. for TIF This Is N₂ Ij = 8/1200 4/120 - T - V, - (1+j) I. -Is = 0 + 258 [165° +82 145°x8/120 + Is = 0 Is =-12.52.1165° +852

please give me a review or comment to that i can help in upcoming question in better way